OpenStudy (anonymous):

3. What is a polynomial function in standard form with zeroes 1, 2, –3, and –3? (1 point)g(x) = x4 + 3x3 –7x2 – 15x + 18 g(x) = x4 + 3x3 –7x2 + 2x + 18 g(x) = x4 – 3x3 + 7x2 + 15x + 18 g(x) = x4 – 3x3 –7x2 + 15x + 18 i think its b

OpenStudy (tamtoan):

why don't you write that out and multiply them and see what you get ? ...if equation have the root of a ...then you can write x - a = 0, use this concept to write out your g(x)

OpenStudy (anonymous):

\[(x-1)(x-2)(x+3)(x-3)=(x-1)(x-2)(x^2-9)\] is a start then more multiplication is needed

OpenStudy (anonymous):

im tring to see if i am right that why i said its is b so am i right yes or no

OpenStudy (anonymous):

i wrote it out wrong. does it really say -3 twice?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

am i right

OpenStudy (anonymous):

it is this one \[x^4+3 x^3-7 x^2-15 x+18\]

OpenStudy (anonymous):

can u show me ach step u did